Sarah is four more than four times as old as Denise. In 8 years the sum of their ages will be 85. How old is Sarah now?
Answer by Paul(988)

Sarah = 4+4x
Denise = x
Sarah in eight years = 8+4+4x
Denise in eight years = 8+x
Sum = 85

(8+4+4x)+(8+x)=85 Simplfy
5x+20=85
5x=65
x=13

4+4(13)=56

Hence, Sarah is 56 years old and Denise is 13 years old.
Paul.